### v1

@@ Abstract
The parallel orbital-updating approach is an orbital iteration based approach for solving eigenvalue problems when many eigenpairs are required, and has been proven to be very efficient, for instance, in electronic structure calculations. In this paper, based on the investigation of a quasi-orthogonality, we present the numerical analysis of the parallel orbital-updating approach for linear eigenvalue problems, including convergence and error estimates of the numerical approximations.
@@ 1 Introduction
Eigenvalue problems are typical models in scientific and engineering computing. For instance, Hartree–Fock type equations and Kohn-Sham equations are widely used mathematical models in electronic structure calculations. The eigenvalues and their corresponding eigenfunctions of these equations provide detailed information about the properties of atoms, molecules, and solids, helping to predict chemical reactions, material properties, and physical behaviors (see e.g. [8, 15, 18, 20]).
@@
In electronic structure calculations of a large system, the approximations of a number of eigenpairs are required. With discretization and the self-consistent field iteration [19, 20, 24], solving the Hartree–Fock type equations or the Kohn-Sham equations is then transformed into repeatedly solving some large scale algebraic eigenvalue problems. It is known that the computational cost of solving such large scale eigenvalue problems is huge. In particular, the solving process often requires large scale orthogonalizing operations, which demand global summation operations and limit the large scale parallelization. Nowadays, the computational scale is limited for systems with hundreds to thousands of atoms. Since applications demand and the supercomputers are available, it is significant to develop scalable and parallelizable numerical methods to solve such eigenvalue problems.
@@
To reduce the computational cost and improve the parallel scalability, a so-called parallel orbital-updating (ParO) approach has been proposed in [9] and developed in [11, 21, 22] for solving eigenvalue problems or their equivalent models resulting from electronic structure calculations. With the ParO approach, we avoid solving the large scale eigenvalue problem and instead solve some independent large scale source problems and some small scale projected eigenvalue problems. Moreover, we see from the numerical experiments in [9, 22] that the stiff matrix resulting from the small scale eigenvalue problem is almost diagonal, which may further reduce the computation cost. Because of the independence, these source problems can be solved in parallel intrinsically. For each source problem, the standard parallel strategies can be applied. It then allows a two-level parallelization: one level of parallelization is obtained by partitioning these source problems into different groups of processors, another level of parallelization is obtained by assigning each source problem to several processors contained in each group. This two-level parallelization demonstrates that the ParO approach has a great potential for large-scale calculations. In fact, the numerical experiments in [9, 11, 22] show the effectiveness of the ParO approach. We conclude that the ParO approach is a powerful parallel computing approach to solving eigenvalue problems, in which many eigenpairs are required. However, up to now, there is no any mathematical justification for the ParO approach.
@@
The purpose of this paper is to present the numerical analysis of the ParO approach for linear eigenvalue problems. We see that, orthogonalizing operations in the process of solving eigenvalue problems, for which the computational cost is very expensive, is usually demanded in the scientific and engineering computing such as electronic structure calculations and quite effects on the efficiency and stability of algorithms [20, 29]. We observe that during the implementation process of the ParO approach, we are able to obtain approximately orthogonal orbitals, which we call quasi-orthogonal orbitals. The ParO approach can be viewed as to utilize the quasi-orthogonal approximations, for which the computational cost is less expensive, to obtain orthogonal approximations. Our numerical analysis is starting from the introduction and investigation of a quasi-orthogonality, which plays a crucial role in the orthogonalization of approximations of the eigenvalue problem. We understand that the presence of both single eigenvalues and multiple eigenvalues renders traditional methods for analyzing single eigenvalues no longer applicable. The difficulty for the case of multiple eigenvalues lies in the fact that the traditional measure for the eigenfunction errors is not valid anymore, because the approximate eigenfunctions obtained in iterations may not approximate the same eigenfunction. Instead of focusing on particular eigenfunctions, in our analysis, we employ the eigenspaces and the gap between the eigenspaces, which brings additional analysing complexities and requires sophisticate functional analysis.
@@
Some approaches for constructing source problems in the ParO approach have been proposed in [9]. As a practical example, the shifted-inverse based ParO algorithm applies the shifted-inverse approach to construct some source problems and solves a small scale eigenvalue problem in each iteration to update the shift parameters to speed up the convergence [9, 22]. To analyze the convergence of the algorithm, we first study its simplified version, which fixes the shift parameters and does not carry out the steps of solving small scale eigenvalue problems in iterations. Under the framework of the ParO approach, we show the convergence of numerical solutions produced by the simplified algorithm, which does not require sufficiently accurate initial guesses. Based on the numerical analysis of the simplified version, we then present a more general and informative convergence result of the shifted-inverse based ParO algorithm than the classical results of the shifted-inverse approach for simple eigenvalues mentioned in, e.g., [2, 23]. To improve the numerical stability, a modified version is proposed in [22], which augments the projected subspace by using the residuals. We also provide a brief outline of the proof for the convergence of this modified algorithm.
@@
The rest of this paper is organized as follows. We recall some existing results of a model problem and introduce the relevant notations in Section 2, and provide some elementary analysis for the quasi-orthogonality in Section 3. In Section 4, we carry out numerical analysis for the ParO approach and its several practical versions. Finally, we give some concluding remarks in Section 5.
@@ 2 Preliminaries
Define $(i,j)<(r,s)$ if $i<r$ or $i=r,j<s$ with $1\leqslant j\leqslant d_{i}$ and $1\leqslant s\leqslant d_{r}$. Let $M(\lambda_{i})$ denote the eigenspace corresponding to $\lambda_{i}$ and $\{u_{ij}\}_{j=1}^{d_{i}}$ be the orthonormal basis of $M(\lambda_{i})$, that is, $M(\lambda_{i})=\operatorname{span}\{u_{i1},\ldots,u_{id_{i}}\}$ for $i=1,2,\ldots$ with $b(u_{ij},u_{kl})=\delta_{ik}\delta_{jl}$, where $\delta_{ik}$ and $\delta_{jl}$ are the Kronecker delta.
@@
A typical example of Eq. 1 is an eigenvalue problem of a partial differential operator over a bounded domain. Let $\Omega\subset\mathbb{R}^{d}(d\geqslant 1)$ be a polygonal domain. We shall use the standard notation for Sobolev spaces $H^{1}(\Omega)$ with associated norms (see, e.g. [1]). Let $H=H_{0}^{1}(\Omega)=\{v\in H^{1}(\Omega):v|_{\partial\Omega}=0\}$ and (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) be the standard $L^{2}$ inner product. Consider the eigenvalue problem: find $\lambda\in\mathbb{R}$ and $u\in H_{0}^{1}(\Omega)$ with $\|u\|_{L^{2}(\Omega)}=1$ such that
@@
We mention that the results obtained in this paper are also valid for a more general bilinear form $a(\cdot,\cdot)$ that
$\|v\|^{2}_{H^{1}_{0}(\Omega)}-C_{1}^{-1}\|v\|_{L^{2}(\Omega)}\leqslant C_{2}a(v,v),\quad\forall w\in H^{1}_{0}(\Omega)$
holds for some constant $C_{1},C_{2}>0$ (see, e.g., Remark 2.9 in [12]).
To carry out the analysis, we apply the following distance between two subspaces $U,V\subset H$ ([7, 14, 17]).
@@
It should be noted that $\operatorname{dist}(u,v)$ is actually the sine of the angle between $u$ and $v$, and independent of the norms of vectors.
Note that $\operatorname{dist}(U,V)=1$ when $\operatorname{dim}(U)>\operatorname{dim}(V)$. We also see that for $U,V,W\subset H$ with $\operatorname{dim}(U)=\operatorname{dim}(V)=\operatorname{dim}(W)<\infty$, there holds that
(2) $\operatorname{dist}(U,V)\leqslant\operatorname{dist}(U,W)+\operatorname{dist}(W,V).$
The following useful lemma can be found in [14, 17].
Lemma 2.2.
Given subspaces $U,V\subset H$, if $\operatorname{dim}(U)=\operatorname{dim}(V)<\infty$, then
$\operatorname{dist}(U,V)=\operatorname{dist}(V,U).$
[…]
Let $V^{h}$ be a finite dimensional subspace of $H$ with $\operatorname{dim}(V^{h})=N_{g}$. The standard finite dimensional discretization of Eq. 1 is defined as follows: find $\lambda^{h}\in\mathbb{R}$ and $0\neq u^{h}\in V^{h}$ such that
@@
We may order the eigenvalues of Eq. 3 as follows:
@@
We assume $\sum_{i=1}^{p}d_{i}=N_{g}(p\geqslant q,i.e.,N_{g}\geqslant N)$. Indeed, the conclusions in this paper hold for all $N_{g}\geqslant N$. The assumption is adopted to simplify the notations in our numerical analysis. Assume that the corresponding eigenfunctions $u^{h}_{ij}$ for $(i,j)\leqslant(p,d_{p})$ satisfy that $b(u^{h}_{ij},u^{h}_{kl})=\delta_{ik}\delta_{jl}$. For $i=1,\ldots,p$, set $M_{h}(\lambda_{i})=\operatorname{span}\{u^{h}_{i1},\ldots,u^{h}_{id_{i}}\}$.
@@
$0\leqslant\lambda^{h}_{ij}-\lambda_{i}\leqslant\lambda^{h}_{ij}\operatorname{dist}_{a}^{2}(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}),\quad\forall(1,1)\leqslant(i,j)\leqslant(q,d_{q}).$
@@
If $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\ll 1$, then there exists $\hat{u}_{ij}\in M(\lambda_{i})$ such that
$\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|_{a}\leqslant L\operatorname{dist}_{a}(M(\lambda_{i}),V^{h}),\quad\forall(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $L$ is a constant that is independent of $V^{h}$.
Note that Proposition 2.4 tells only that each orthonormal eigenfunction of Eq. 3 approximates some eigenfunction of Eq. 1, which may not be orthonormal each other (see Corollary 2.11 in [10]). However, in practical applications, the approximate property between the orthonormal approximate eigenfunctions and the orthonormal exact eigenfunctions are usually required, which are of structure-preserving and can be used to prevent the accumulation of errors of approximations.
@@ 3 Quasi-orthogonality
To carry out the numerical analysis for the ParO approach, we introduce and investigate a quasi-orthogonality, which plays a crucial role in the orthogonalization of approximations of the eigenvalue problem.
Let $\{v_{j}\}_{j=1}^{n}\subset H$ be linearly independent. Consider the Gram-Schmidt orthogonalization of $\{v_{j}\}_{j=1}^{n}$:
  • Choose $\tilde{v}_{1}=v_{1}.$
  • For $j=2,3,\ldots,n,$ set
    (4) $\tilde{v}_{j}=v_{j}-\sum_{l=1}^{j-1}\frac{a(\tilde{v}_{l},v_{j})}{\left\|\tilde{v}_{l}\right\|_{a}^{2}}\tilde{v}_{l}.$
We have the following useful lemma, which tells the properties of the orthogonalization of quasi-orthogonal vectors.
Lemma 3.1.
Given $\theta\in(0,1)$ and $\delta\in\left(0,\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\right)$, let $\{u_{j}\}_{j=1}^{n},\{v_{j}\}_{j=1}^{n}\subset H$ satisfy
(5) $a(u_{i},u_{j})=\delta_{ij},\,\|v_{j}\|_{a}=1,\quad\|u_{j}-v_{j}\|_{a}\leqslant\delta,\quad i,j=1,2,\ldots,n,$
where $\{v_{j}\}^{n}_{j=1}$ is said to be quasi-orthogonal. If $\{\tilde{v}_{j}\}_{j=1}^{n}$ is obtained by the Gram-Schmidt orthogonalization of $\{v_{j}\}_{j=1}^{n}$, then for $j=2,3,\ldots,n$, there holds
$\left\|\tilde{v}_{j}-v_{j}\right\|_{a}\leqslant\frac{1}{\theta}M_{j}\delta,$
where $M_{j}=\frac{2}{\theta^{j-2}}\left((1+\theta)^{j-1}-\theta^{j-1}\right)$.
Proof 3.2.
To obtain the conclusion, we see from the Eq. 4 that it is sufficient to prove
(6) $\sum_{l=1}^{j-1}|a(\tilde{v}_{l},v_{j})|\leqslant M_{j}\delta,\quad\|\tilde{v}_{j}\|_{a}\geqslant\theta,$
for any $j=2,3,\ldots,n$. We prove Eq. 6 by induction.
Note that
$a(v_{1},v_{2})=a(v_{1}-u_{1},v_{2})+a(u_{1},v_{2}-u_{2}),$
which implies
$|a(v_{1},v_{2})|\leqslant\|v_{1}-u_{1}\|_{a}+\|v_{2}-u_{2}\|_{a}\leqslant 2\delta.$
The facts $\tilde{v}_{1}=v_{1},\|v_{1}\|_{a}=\|u_{1}\|_{a}=1$, and
$\tilde{v}_{2}=v_{2}-\frac{a(\tilde{v}_{1},v_{2})}{\|\tilde{v}_{1}\|_{a}^{2}}\tilde{v}_{1}$
yield
$|a(\tilde{v}_{1},v_{2})|\leqslant 2\delta,\quad\|\tilde{v}_{2}\|_{a}\geqslant 1-2\delta,\quad\|\tilde{v}_{2}-v_{2}\|_{a}\leqslant 2\delta.$
A simple calculation shows that $1-2\delta\geqslant\theta$. Hence, Eq. 6 is true when $j=2$.
We assume Eq. 6 is true for $2\leqslant j\leqslant k<n$. We obtain from Eq. 6 that
(7) $\|\tilde{v}_{j}-v_{j}\|_{a}\leqslant\sum_{l=1}^{j-1}\frac{|a(\tilde{v}_{l},v_{j})|}{\|\tilde{v}_{l}\|_{a}}\leqslant\frac{M_{j}\delta}{\theta},\quad j=2,3,\ldots,k.$
Note that Eq. 5 and the identity
$a(\tilde{v}_{l},v_{k+1})=a(\tilde{v}_{l}-v_{l},v_{k+1})+a(v_{l}-u_{l},v_{k+1})+a(u_{l},v_{k+1}-u_{k+1}),\quad l=1,2,\ldots k$
imply
$|a(\tilde{v}_{l},v_{k+1})|$ $\leqslant\|\tilde{v}_{l}-v_{l}\|_{a}+2\delta,\quad l=1,2,\ldots,k.$
Thus we have
$\sum_{l=1}^{k}|a(\tilde{v}_{l},v_{k+1})|\leqslant 2k\delta+\sum_{l=2}^{k}\|\tilde{v}_{l}-v_{l}\|_{a},\quad k\geqslant 2.$
We obtain from Eq. 6 that
$\|\tilde{v}_{k}-v_{k}\|_{a}\leqslant\frac{1}{\theta}\sum_{l=1}^{k-1}|a(\tilde{v}_{l},v_{k})|\leqslant\frac{M_{k}\delta}{\theta}.$
Then we arrive at
$\sum_{l=1}^{k}|a(\tilde{v}_{l},v_{k+1})|\leqslant M_{k+1}\delta,$
where $M_{k+1}=2k+\frac{1}{\theta}\sum_{l=1}^{k}M_{l}$, i.e., $M_{k+1}=\frac{2}{\theta^{k-1}}\left((1+\theta)^{k}-\theta^{k}\right).$
Due to $\tilde{v}_{k+1}=v_{k+1}-\sum_{l=1}^{k}\frac{a(\tilde{v}_{l},v_{k+1})}{\|\tilde{v}_{l}\|_{a}^{2}}\tilde{v}_{l},$ we have
$\|\tilde{v}_{k+1}\|_{a}\geqslant 1-\sum_{l=1}^{k}\frac{|a(\tilde{v}_{l},v_{k+1})|}{\|\tilde{v}_{l}\|_{a}}\geqslant 1-\frac{M_{k+1}\delta}{\theta}.$
Since $M_{k+1}\leqslant\frac{2}{\theta^{n-2}}\left((1+\theta)^{n-1}-\theta^{n-1}\right)$ is true for $2\leqslant k<n$ and
$\delta<\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})},$
we conclude that $\|\tilde{v}_{k+1}\|_{a}\geqslant\theta,$ which proves Eq. 6.
After a simple calculation, we have the following conclusion.
@@
Given $\theta\in(0,1)$ and $\delta\in\left(0,\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\right)$, let $\{u_{j}\}_{j=1}^{n},\{v_{j}\}_{j=1}^{n}\subset H$ satisfy
$a(u_{i},u_{j})=\delta_{ij},\|v_{j}\|_{a}=1\quad i,j=1,2,\ldots,n;$
$\|u_{j}-v_{j}\|_{a}\leqslant\delta,\quad~{}j=1,2,\ldots,n.$
Then there exists $\{w_{j}\}_{j=1}^{n}\subset\operatorname{span}\left\{v_{1},\ldots,v_{n}\right\}$ such that
$a(w_{i},w_{j})$ $=\delta_{ij},\quad i,j=1,2,\ldots,n;$
$\operatorname{dist}_{a}\left(u_{j},w_{j}\right)$ $\leqslant\left(1+\frac{2}{\theta^{n-3}}\left(\left(1+\theta\right)^{n-1}-\theta^{n-1}\right)\right)\delta,\quad j=1,2,\ldots,n.$
For given $\varepsilon\in(0,1)$, we consider $U=\operatorname{span}\{u_{1},\ldots,u_{n}\}\subset H$ satisfying $a(u_{i},u_{j})=\delta_{ij}$ and $V\subset H$ satisfying $\operatorname{dim}(V)=\operatorname{dim}(U)$ and $\operatorname{dist}_{a}(U,V)\leqslant\varepsilon$.
Next we show that $\mathcal{P}_{V}|_{U}$ is an isomorphism from $U$ to $V$. Indeed, for $\tilde{u},\hat{u}\in U$ satisfying $\mathcal{P}_{V}\tilde{u}=\mathcal{P}_{V}\hat{u}$, we obtain from

### v2

@@ Abstract
The parallel orbital-updating approach is an orbital/eigenfunction iteration based approach for solving eigenvalue problems when many eigenpairs are required. It has been proven to be efficient, for instance, in electronic structure calculations. In this paper, based on the investigation of a quasi-orthogonality, we present the numerical analysis of the parallel orbital-updating approach for linear eigenvalue problems, including convergence and error estimates of the numerical approximations.
@@ 1 Introduction
Eigenvalue problems are typical models in scientific and engineering computing. For instance, Hartree–Fock type and Kohn-Sham equations are widely used mathematical models in electronic structure calculations. The eigenvalues and their corresponding eigenfunctions of these equations provide detailed information about the properties of atoms, molecules, and solids, helping to predict chemical reactions, material properties, and physical behaviors (see e.g. [9, 17, 20, 22]).
@@
In electronic structure calculations of a large system, the approximations of many eigenpairs are required. With discretization and the self-consistent field iteration [21, 22, 26], solving the Hartree–Fock type equations or the Kohn-Sham equations is then transformed into repeatedly solving some large scale algebraic eigenvalue problems. It is known that the computational cost of solving such large scale eigenvalue problems is high. In particular, the solving process often requires large scale orthogonalizing operations, which demand global summation operations and limit large scale parallelization. Nowadays, the computational scale is limited for systems with hundreds to thousands of atoms. Since applications demand and supercomputers are available, it is significant to develop scalable and parallelizable numerical methods to solve such eigenvalue problems.
@@
To reduce the computational cost and improve the parallel scalability, a so-called parallel orbital-updating (ParO) approach has been proposed in [10] and developed in [12, 23, 24] for solving eigenvalue problems and their equivalent models resulting from electronic structure calculations. We mention that there are also some other methods for approximating eigenpairs in the literature such as the density matrix based algorithm [27], the subspace iteration algorithm [31], and the projection method based on the root-finding of the analytic function [32]. With ParO, we avoid solving the large scale eigenvalue problem directly and instead solve some independent large scale source problems and small scale projected eigenvalue problems to obtain approximate eigenpairs. Moreover, we see from the numerical experiments in [10, 24] that the stiff matrix corresponding to the small scale eigenvalue problem is almost diagonal, which may further reduce the computation cost. Because of their independence, these source problems can be solved intrinsically in parallel. For each source problem, the standard parallel strategies can be applied. It then allows a two-level parallelization: one level of parallelization is obtained by partitioning these source problems into different groups of processors, and another level of parallelization is obtained by assigning each source problem to several processors contained in each group. This two-level parallelization demonstrates that ParO has great potential for large scale calculations. The numerical experiments in [10, 12, 24] show that ParO is efficient, of good scalability of parallelization, and can produce highly accurate approximations. We conclude that ParO is a powerful parallel computing approach for solving eigenvalue problems, in which many eigenpairs are required, and has been integrated into the electronic structure calculation software Quantum ESPRESSO [1]. However, up to now, there is no mathematical justification for ParO.
@@
The purpose of this paper is to present the numerical analysis of ParO for linear eigenvalue problems. We observe that during the implementation process of ParO, we are able to obtain approximately orthogonal eigenfunctions, which we call quasi-orthogonal eigenfunctions. ParO can be viewed as utilizing the quasi-orthogonal approximations, for which the computational cost is lower, to obtain orthogonal approximations. Our numerical analysis starts from the introduction and investigation of a quasi-orthogonality, which plays a crucial role in our investigations on approximations of eigenvalue problems. We understand that the presence of both single eigenvalues and multiple eigenvalues renders traditional methods for analyzing single eigenvalues no longer applicable. The difficulty for the case of multiple eigenvalues lies in the fact that the traditional measure for the eigenfunction errors is not valid anymore, because the approximate eigenfunctions obtained in iterations may not approximate the same eigenfunction. Instead of focusing on particular eigenfunctions, in our analysis, we employ the eigenspaces and the gap from the eigenspaces to their approximations, which brings additional analyzing complexities and requires sophisticated functional analysis.
@@
Some approaches for constructing source problems in ParO have been proposed in [10]. As a practical example, the shifted-inverse based ParO algorithm applies the shifted-inverse approach to construct some source problems and solves a small scale eigenvalue problem in each iteration to update the shifts to speed up the convergence [10, 24]. To analyze the convergence of the algorithm, we first study its simplified version, which fixes the shifts and does not carry out the steps of solving small scale eigenvalue problems in iterations. Within the framework of ParO, we demonstrate the convergence of numerical solutions produced by the simplified algorithm, which does not require sufficiently accurate initial guesses. Based on the numerical analysis of the simplified version, we then present a more general and informative convergence result of the shifted-inverse based ParO algorithm than the classical results of the shifted-inverse approach for simple eigenvalues mentioned in, e.g., [3, 25]. To improve the numerical stability, a modified version is proposed in [24], which augments the projected subspace by using the residuals. We also provide a brief outline of the proof for the convergence of this modified algorithm.
@@
The rest of this paper is organized as follows. We recall some existing results of a model problem and introduce the relevant notation in Section 2, and provide some elementary analysis for the quasi-orthogonality in Section 3. In Section 4, we show the convergence and error estimates of the numerical approximations produced by ParO and its practical versions. We give some concluding remarks in Section 5. Finally, we present the corresponding detailed proofs in Appendix A.
@@ 2 Preliminaries
For $1\leqslant j\leqslant d_{i}$ and $1\leqslant s\leqslant d_{r}$, denote $(i,j)<(r,s)$ if $i<r$ or $i=r,j<s$. Let $M(\lambda_{i})$ be the eigenspace corresponding to $\lambda_{i}$ and $\{u_{ij}\}_{j=1}^{d_{i}}$ be the orthonormal basis of $M(\lambda_{i})$, that is, $M(\lambda_{i})=\operatorname{span}\{u_{i1},\ldots,u_{id_{i}}\}$ for $i=1,2,\ldots$ with $b(u_{ij},u_{kl})=\delta_{ik}\delta_{jl}$, where $\delta_{ik}$ and $\delta_{jl}$ are the Kronecker delta.
@@
An example
A typical example of Eq. 1 is an eigenvalue problem of a partial differential operator over a bounded domain. Let $\Omega\subset\mathbb{R}^{d}(d\geqslant 1)$ be a bounded domain. We shall use the standard notation for Sobolev spaces $H^{1}(\Omega)$ with associated norms (see, e.g., [2]). Let $H=H_{0}^{1}(\Omega)=\{v\in H^{1}(\Omega):v|_{\partial\Omega}=0\}$ and (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) be the standard $L^{2}$ inner product. Consider the eigenvalue problem: find $\lambda\in\mathbb{R}$ and $u\in H_{0}^{1}(\Omega)$ with $\|u\|_{L^{2}(\Omega)}=1$ such that
@@
We mention that the results obtained in this paper are also valid for a more general bilinear form $a(\cdot,\cdot)$ with
$\|w\|^{2}_{H^{1}_{0}(\Omega)}-C_{1}^{-1}\|w\|_{L^{2}(\Omega)}\leqslant C_{2}a(w,w),\quad\forall w\in H^{1}_{0}(\Omega)$
holding for some constant $C_{1},C_{2}>0$ (see, e.g., Remark 2.9 in [13]).
Distance from one subspace to another
To carry out the analysis, we apply the following distance from the nontrivial subspace $U\subset H$ to the subspace $V\subset H$ ([8, 16, 19]):
@@
We see that $\operatorname{dist}(u,v)$ is actually the sine of the angle between $u$ and $v$, and independent of the norms of vectors. Note that $\operatorname{dist}(U,V)=1$ when $\operatorname{dim}(U)>\operatorname{dim}(V)$. We also observe that for $U,V\subset H$ with $\operatorname{dim}(U)=\operatorname{dim}(V)<\infty$, there holds that
(2) $\operatorname{dist}(U,V)=\operatorname{dist}(V,U).$
Define $\mathcal{P}_{V}$ to be the orthogonal projection from $H$ onto $V\subset H$ with respect to the inner product $a(\cdot,\cdot)$.
Finite dimensional approximation
Let $V^{h}$ be a finite dimensional subspace of $H$ with $\operatorname{dim}(V^{h})=N_{g}\geqslant N$. The standard finite dimensional discretization of Eq. 1 is defined as follows: find $\lambda^{h}\in\mathbb{R}$ and $0\neq u^{h}\in V^{h}$ such that
@@
For convenience, we assume that there exists $p\geqslant q$ such that $N_{g}=\sum_{i=1}^{p}d_{i}$. Then we may order the eigenvalues of Eq. 3 as follows:
@@
The assumption is adopted to simplify the notation in our numerical analysis. In fact, our analysis results in this paper hold for all $N_{g}\geqslant N$, regardless of whether $N_{g}=\sum_{i=1}^{p}d_{i}$ is satisfied. Assume that the corresponding eigenfunctions $u^{h}_{ij}$ for $(i,j)\leqslant(p,d_{p})$ satisfy that $b(u^{h}_{ij},u^{h}_{kl})=\delta_{ik}\delta_{jl}$. For $i=1,\ldots,p$, set $M_{h}(\lambda_{i})=\operatorname{span}\{u^{h}_{i1},\ldots,u^{h}_{id_{i}}\}$.
@@
$0\leqslant\lambda^{h}_{ij}-\lambda_{i}\leqslant\lambda^{h}_{ij}\operatorname{dist}^{2}(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}),\quad\forall(1,1)\leqslant(i,j)\leqslant(q,d_{q}).$
@@
Let $\{u_{ij}^{h}\}$ be the solutions of Eq. 3. There exists $\epsilon_{*}\in(0,1)$ satisfying that if $\operatorname{dist}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\leqslant\epsilon_{*}$, there exists $\hat{u}_{ij}\in M(\lambda_{i})$ such that
$\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|\leqslant C_{*}\operatorname{dist}(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}),\quad\forall(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $C_{*}$ is a constant that is independent of $V^{h}$.
Remark 2.4.
Based on Theorem 1 and Theorem 2 in [18], the constants in Proposition 2.3 can be chosen as
$\epsilon_{*}=\alpha\min_{i=1,\ldots,q+1}\sqrt{\frac{\lambda_{i}-\lambda_{i-1}}{\lambda_{i}}}\quad\text{and}\quad C_{*}=\max_{i=1,\ldots,q}\sqrt{\frac{2(\lambda_{i+1}-\lambda_{i-1})\lambda_{i}}{(1-\alpha^{2})(\lambda_{i}-\lambda_{i-1})(\lambda_{i+1}-\lambda_{i})}}$
with $\alpha\in(0,1)$.
@@ 3 Quasi-orthogonality
To understand the philosophy behind ParO and carry out the numerical analysis, we introduce a quasi-orthogonality, which plays a crucial role in our investigations on approximations of eigenvalue problems.
Definition 3.1.
Given $\delta>0$ and $n\geqslant 2$, $\{v_{j}\}^{n}_{j=1}\subset H$ with $\|v_{j}\|=1$ for $j=1,\ldots,n$ is said to be $\delta$-quasi-orthogonal if there exists $\{u_{j}\}^{n}_{j=1}\subset H$ satisfying that
(4) $a(u_{i},u_{j})=\delta_{ij},,\quad\|u_{j}-v_{j}\|\leqslant\delta,\quad i,j=1,2,\ldots,n.$
The following proposition tells the approximation property of the orthogonalization of quasi-orthogonal vectors. The proof is provided in Section A.1.
Proposition 3.2.
If $\{v_{j}\}_{j=1}^{n}\subset H$ is $\delta$-quasi-orthogonal, then there exists an orthonormal basis $\{\tilde{v}_{j}\}_{j=1}^{n}$ of $\operatorname{span}\{v_{1},\ldots,v_{n}\}$ such that
$\left\|\tilde{v}_{j}-v_{j}\right\|\leqslant\sqrt{n}\delta,\quad j=1,\ldots,n.$
The following conclusion, which can be derived directly by a triangle inequality from Proposition 3.2, shows the approximation property between orthonormal bases.
@@
If $\{v_{j}\}_{j=1}^{n}\subset H$ is $\delta$-quasi-orthogonal, i.e., there exists $\{u_{j}\}_{j=1}^{n}\subset H$ satisfying
$a(u_{i},u_{j})=\delta_{ij},\quad\|u_{j}-v_{j}\|\leqslant\delta,\quad~{}i,j=1,2,\ldots,n,$
then there exists $\{\tilde{v}_{j}\}_{j=1}^{n}\subset\operatorname{span}\{v_{1},\ldots,v_{n}\}$ such that
$a(\tilde{v}_{i},\tilde{v}_{j})$ $=\delta_{ij},\quad\operatorname{dist}(u_{j},\tilde{v}_{j})\leqslant\left\|u_{j}-\tilde{v}_{j}\right\|\leqslant(\sqrt{n}+1)\delta,\quad i,j=1,\ldots,n.$
Orthogonal basis approximation
We arrive at the following proposition from Corollary 3.3, which tells the approximation property of orthonormal bases by the distance from one subspace to another and will play a crucial role in our analysis.
Proposition 3.4.
Given $\varepsilon\in(0,1)$ and two subspaces $U,V\subset H$ with $\operatorname{dim}(U)=\operatorname{dim}(V)=n$. If $\operatorname{dist}(U,V)\leqslant\varepsilon,$ then for any orthonormal basis $\{u_{j}\}_{j=1}^{n}$ of $U$, there exists an orthonormal basis $\{w_{j}\}_{j=1}^{n}$ of $V$ satisfying
$\operatorname{dist}(u_{i},w_{i})\leqslant(1+\sqrt{n})\sqrt{2-2\sqrt{1-\varepsilon^{2}}},\quad i,j=1,2,\ldots,n,$
Proof 3.5.
First, we show that $\mathcal{P}_{V}|_{U}$ is an isomorphism from $U$ to $V$. Indeed, for $\tilde{u},\hat{u}\in U$ satisfying $\mathcal{P}_{V}\tilde{u}=\mathcal{P}_{V}\hat{u}$, we obtain from